OpenStudy (anonymous):

In a 30-60-90 triangle, the short leg is 8 m. Find the length of the long leg. Answer 16 m 11.31 m 4.62 m 13.86 m

OpenStudy (angela210793):

16 cause the longest leg is 2*shortest leg=8*2=16 (the leg in front of 30 degree angle is hypotenuse/2) so hypotenuse=2*shortest leg

OpenStudy (anonymous):

thats not right..

OpenStudy (anonymous):

thats not right

OpenStudy (amistre64):

the legs of a 30-60 are: 1 and sqrt(3) since the shortest one is 8; multiply the other by 8 to get 8sqrt(3)

OpenStudy (angela210793):

O.o...ok just ignore my answer.....:/

OpenStudy (amistre64):

i believe the logic goes; you didnt give me the answer in exact form and i dont want to waste my precious time in figuring out the answer. i could be wrong :)

OpenStudy (angela210793):

i think i'm right.....(or maybe i forgot math) I miss school :/